Grade 12Mechanics

You drive on Interstate 10 from San Antonio to Houston, half the time at 55 km/h and the other half at 90 km/h. On the way back you travel half the distance at 55 km/h and the other half at 90 km/h. What is your average speed (a) from San Antonio to Houston, (b) from Houston back to San Antonio, and (c) for the entire trip? (d) What is your average velocity for the entire trip? (e) Sketch x versus t for (a), assuming the motion is all in the positive x direction. Indicate how the average velocity can be found on the sketch.

Request : I have Solved Part (a) , (b),(c) and (d) respectively. I am confused with the graph given in part (e) please give the answer of graph part.

Thanks !

Ashish Ranjan
10 Years agoGrade 12

1 Answer

Vikas TU
10 Years ago
a) avg. speed = (total distance)/(total time) = (55*t/2 + 90*t/2)/t = (90 + 50)/2 = 70 km/hr.
  1. avg. speed will  remain same.
  2. for entire trip,
---------> it will also remain same.
  1. avg. velocity = total displacement / total time = 0 for entire trip.